# Transport, Variational Inference and Diffusions: with Applications to Annealed Flows and Schrödinger Bridges

### v1

@@ 3.1 Score-based annealed flows
Consider the log-variance loss, which can be re-expressed as
$\mathcal{L}(\phi):=\mathrm{Var}\left(\operatorname{ln}\frac{\mathrm{d}\overrightarrow{{\mathbb{P}}}^{\pi_{0},\sigma^{2}\nabla\operatorname{ln}\pi+\nabla\phi}}{\mathrm{d}\overleftarrow{{\mathbb{P}}}^{\pi_{T},-\sigma^{2}\nabla\operatorname{ln}\pi+\nabla\phi}}\right)=$
$\!\!\!=\mathrm{Var}\Bigg{(}\operatorname{ln}\pi_{T}({\bm{Y}}_{T})-\operatorname{ln}\pi_{0}({\bm{Y}}_{0})+\int_{0}^{T}\Delta\phi_{t}({\bm{Y}}_{t})\,\mathrm{d}t$
$\!\!\!\!-\!\!\sigma\sqrt{2}\!\!\!\int_{0}^{T}\!\!\!\!\!\nabla\operatorname{ln}\pi_{t}({\bm{Y}}_{t})\!\circ\!\mathrm{d}{\bm{W}}_{t}\!-\sigma^{2}\!\!\int_{0}^{T}\!\!\!|\nabla\operatorname{ln}\pi_{t}({\bm{Y}}_{t})|^{2}\,\!\mathrm{d}t\!\!\Bigg{)},$ (19)
see Appendix D. Here, $\circ\mathrm{d}{\bm{W}}_{t}$ denotes Stratonovich integration, and the variance is taken with respect to samples from (3.1). We note that in the limit $\sigma\rightarrow 0$, the objective in (4) enforces an integrated version of the instantaneous change of density formula $\partial_{t}\operatorname{ln}\pi_{t}({\bm{Y}}_{t})=-\Delta\phi_{t}({\bm{Y}}_{t})$ for continuous time normalising flows of the form $\dot{{\bm{Y}}_{t}}=\nabla\phi_{t}({\bm{Y}}_{t})$, see Papamakarios et al. (2021, Section 4). Notwithstanding the intuitive appeal of the term involving $\Delta\phi_{t}$, it may be favourable computationally to replace it by a combination of forward and backward integrals using (16), see Remark 13.
The additional constraints detailed above resolve the nonuniqueness in Frameworks 1 and ${}^{\prime}$:
Proposition 3.1.
Under mild conditions on $(\pi_{t})_{t\in[0,T]}$, (18) admits a ($\pi_{t}$-a.e.)unique minimiser, up to additive constants.
We provide a proof in Appendix D; a numerical evaluation is postponed to future work.
Remark 5 (Detaching the gradient).
Since the log-variance divergence is valid regardless of the base distribution according to which the variance is computed, it is possible to detach the gradient with respect to ${\bm{Y}}_{t}$ when optimising (4) (Nüsken and Richter, 2021; Richter et al., 2020), avoiding differentiation through the score $\nabla\operatorname{ln}\pi_{t}$ from which a $D_{\mathrm{KL}}$-based objective would suffer.
Remark 6 (Action matching).
Neklyudov et al. similarly propose to learn $\phi_{t}$ in such a way that (3.1) reproduces a fixed curve of distributions $(\pi_{t})_{t\in[0,T]}$, albeit from samples rather than from available scores. We show in Appendix C.4 that their objective can be recovered from Framework ${}^{\prime}$ using a diffusion with marginals $(\pi_{t})_{t\in[0,T]}$ as the reference process.
Finally following Zhang and Chen (2022); Vargas et al. (2023a), we can obtain an unbiased estimate of the normalising $Z$ in $\pi_{T}=\hat{\pi}_{T}/Z$ from the RND in (4) via the following importance sampling identity:
$S({\bm{Y}})\!=\operatorname{ln}\pi_{T}({\bm{Y}}_{T})-\operatorname{ln}$ $\pi_{0}({\bm{Y}}_{0})+\int_{0}^{T}\Delta\phi_{t}({\bm{Y}}_{t})\,\mathrm{d}t$
$\hat{Z}$ $=e^{-S({\bm{Y}})},$ (3.1)
where, as mentioned previously, we can use the more computationally friendly expression for the RND in equation (F.1), which as in Vargas et al. (2023a) provides an ELBO for $\hat{Z}$ when discretised (see Remark 17).
3.2 Fluctuation theorem and the Jarzynski equality
To shed further light on score-based annealed flows, we examine the RND involved in (4) when $\phi=0$. Connecting (3.1) to well-established importance sampling (IS) estimators for partition functions (Jarzynski, 1997; Neal, 2001; Chopin, 2002) we will motivate how our approach in Section 3.1 allows us to improve on said estimators. At the same time, we showcase an application of Proposition 2.2.
To establish the aforementioned connection we present a novel derivation of the Jarzynski equality (Jarzynski, 1997) and Crooks’ fluctuation theorem (Crooks, 1999) for diffusion processes which is more in line with Crooks’ original proof. Our derivation appears to be more succinct and ‘pathwise’, in contrast to previous approaches which combine the Feynman-Kac formula with the FPK equation (Chen et al., 2019; Stoltz et al., 2010; Hartmann et al., 2019; Hummer and Szabo, 2001).
@@ 3.2 Fluctuation theorem and the Jarzynski equality
Proposition 3.2 (Crooks’ fluctuation theorem).
Given the annealed flow $\pi_{t}=\hat{\pi}_{t}/{\mathcal{Z}}_{t}$ interpolating between $\pi_{0}$ and $\pi_{T}$, we define work done and free energy as
${\mathcal{W}}_{T}({\bm{Y}})=-\int_{0}^{T}\sigma^{2}\partial_{t}\ln\hat{\pi}_{t}({\bm{Y}}_{t})\,\mathrm{d}t,\quad{\mathcal{F}}_{t}=-{\sigma^{2}}\ln{\mathcal{Z}}_{t},$
see Jarzynski (1997); Chen et al. (2019). Then we have Crooks’ identity,
@@
$\left(\frac{\mathrm{d}\overrightarrow{\mathbb{P}}^{\pi_{0},\sigma^{2}\nabla\ln{\pi}}}{\mathrm{d}\overleftarrow{\mathbb{P}}^{\pi_{T},-\sigma^{2}\nabla\ln{\pi}}}\right)({\bm{Y}})=e^{-\frac{1}{\sigma^{2}}({\mathcal{F}}_{T}-{\mathcal{F}}_{0})}e^{\frac{1}{\sigma^{2}}{\mathcal{W}}_{T}({\bm{Y}})},$
$\overrightarrow{\mathbb{P}}^{\pi_{0},\sigma^{2}\nabla\ln{\pi}}$-a.s., which implies Jarzynski’s equality
@@
$\mathbb{E}_{\overrightarrow{\mathbb{P}}^{\pi_{0},\sigma^{2}\nabla\ln\pi}}\left[e^{-\frac{1}{\sigma^{2}}{\mathcal{W}}_{T}}\right]=\frac{{\mathcal{Z}}_{T}}{{\mathcal{Z}}_{0}}$ (21)
by taking expectations.
The proof directly uses Proposition 2.2 and the conversion formula (16) to compute the RND ${\mathrm{d}\overrightarrow{\mathbb{P}}^{\pi_{0},\sigma^{2}\nabla\ln{\pi}}}/{\mathrm{d}\overleftarrow{\mathbb{P}}^{\pi_{T},-\sigma^{2}\nabla\ln{\pi}}}$, followed by an application of Itô’s formula to $t\mapsto\operatorname{ln}\hat{\pi}_{t}({\bm{Y}}_{t})$, see Appendix E.2.
Prior works (Vaikuntanathan and Jarzynski, 2008; Neal, 2001; Chopin, 2002) have made use of the Jarzynski equality (21) to estimate partition functions via importance sampling. However, such choices for forward and backward processes do not yield optimal variance for the estimator (Del Moral et al., 2006), and indeed standard Monte Carlo estimators for (21) might exhibit high variance, see Stoltz et al. (2010, Section 4.1.4). In contrast, our proposed approach allows us to optimise the vector field $\nabla\phi_{t}$ which if optimised fully would make the IS estimator (3.1) optimal (zero variance).
Remark 7 (Related work).
The task of learning the vector field $\nabla\phi_{t}$ so that (3.1) reproduces $(\pi_{t})_{t\in[0,T]}$ has been approached from various directions. Reich (2011); Heng et al. (2021); Reich (2022); Vaikuntanathan and Jarzynski (2008) explore methodologies that exploit the characterisation of $\nabla\phi_{t}$ in terms of the elliptic PDE (55). Arbel et al. (2021) propose to leverage normalising flows sequentially to minimise KL divergences between implied neighboring densities. In an appropriate limiting regime, they recover the SDE (3.1), see Remark 14. More closely related to the objective in (4) are physics-inspired attempts to optimise the IS-variance in (21), see (Hartmann et al., 2019; Doucet et al., 2022; Zhang, 2021), however, these methods are based on the time reversal of ULA diffusions and unlike ours do not reproduce the interpolating distribution $(\pi_{t})_{t\in[0,T]}$.
4 Entropic optimal transport
Another way of selecting a particular transition between $\mu$ and $\nu$ is by imposing an entropic penalty, encouraging the dynamics to stay close to a prescribed, oftentimes physically or biologically motivated, reference process. Using the notation employed in Framework 1, the static Schrödinger problem (Schrödinger, 1931) is given by
$\pi^{*}({\bm{x}},{\bm{z}})\in\operatorname*{arg\,min}_{\pi({\bm{x}},{\bm{z}})}\Big{\{}D_{\mathrm{KL}}(\pi({\bm{x}},{\bm{z}})||r({\bm{x}},{\bm{z}})):$
$\pi_{\bm{x}}({\bm{x}})=\mu({\bm{x}}),\pi_{\bm{z}}({\bm{z}})=\nu({\bm{z}})\Big{\}},$ (22)
where $r({\bm{x}},{\bm{z}})$ is the Schrödinger prior encoding additional domain-specific information. In the context of Framework ${}^{\prime}$, the dynamical Schrödinger problem can be presented in the form
@@ 4 Entropic optimal transport
*${\mathbb{P}}^{*}\!\!\in\!\operatorname*{arg\,min}_{\overrightarrow{{\mathbb{P}}}^{\mu,a}_{T}\;=\;\nu}\mathbb{E}_{{\bm{Y}}\sim\overrightarrow{{\mathbb{P}}}^{\mu,a}}\!\!\left[\tfrac{1}{2\sigma^{2}}\!\int_{0}^{T}\!\!\|a_{t}-f_{t}\|^{2}({\bm{Y}}_{t})\,\mathrm{d}t\right]\!,$ (23)
that is, the driving vector field $a_{t}$ determining ${\mathbb{P}}^{*}$ should be chosen in such a way that (i), the corresponding diffusion transitions from $\mu$ to $\nu$, and (ii), among such diffusions, the vector field $a_{t}$ remains close to the prescribed vector field $f_{t}$, in mean square sense. Under mild conditions, it can be shown that solutions to the problems in (4) and (23) exist and are unique, and that the static and dynamic viewpoints are related through a mixture-of-bridges construction (assuming that the conditionals $r({\bm{z}}|{\bm{x}})$ correspond to the transitions induced by $f_{t}$), [^5] This correspondence shows that the solution in (4) only depends on the conditionals $r({\bm{z}}|{\bm{x}})$, and not on the full distribution $r({\bm{x}},{\bm{z}})$. see (Léonard, 2014a, Section 2). To make this relationship plausible, we remark that the $D_{\mathrm{KL}}$-divergence in (4) can be translated into the time integral in (23) using Girsanov’s theorem.
Table 1: Generative Modelling Results comparing DNF (Zhang and Chen, 2021) ($\lambda=0$) to our PINN regualirsed approach with $\lambda=0.5$. We observe that PINN regularisation obtains similar KL and Cross entropy losses to DNF whilst achieving lower distances to the prior.
(a) $\lambda=2$
(b) $\lambda=0$ (EM Init)
(c) $\lambda=0$ (Random Init)
(d) DNF (No Prior)
Figure 1: (0(a)) our proposed regularised objective, (0(b)) $\lambda$ set to 0 but using clever EM motivated initialisation, (0(c)) $\lambda$ set to 0 with random initialisation of the forward drift, (0(d)) for reference DNF with $f_{t}=0$ (uninformative Schrödinger prior).
4.1 Iterative proportional fitting (IPF) and the EM algorithm
It is well known that approximate solutions for $\pi^{*}({\bm{x}},{\bm{z}})$ and ${\mathbb{P}}^{*}$ can be obtained using alternating $D_{\mathrm{KL}}$-projections, keeping one of the marginals fixed in each iteration: Under mild conditions, the sequence defined by
$\pi^{2n+1}({\bm{x}},{\bm{z}})=$ (24a)
@@ 4.1 Iterative proportional fitting (IPF) and the EM algorithm
$\operatorname*{arg\,min}_{\pi({\bm{x}},{\bm{z}})}\left\{D_{\mathrm{KL}}(\pi({\bm{x}},{\bm{z}})||\pi^{2n}({\bm{x}},{\bm{z}})):\,\,\pi_{\bm{x}}({\bm{x}})=\mu({\bm{x}})\right\},$
$\pi^{2n+2}({\bm{x}},{\bm{z}})=$ (24b)
$\operatorname*{arg\,min}_{\pi({\bm{x}},{\bm{z}})}\left\{D_{\mathrm{KL}}(\pi({\bm{x}},{\bm{z}})||\pi^{2n+1}({\bm{x}},{\bm{z}})):\,\,\pi_{\bm{z}}({\bm{z}})=\nu({\bm{z}})\right\},$
for $n\geq 0$, with initialisation $\pi^{0}({\bm{x}},{\bm{z}})=r({\bm{x}},{\bm{z}})$, is known to converge to $\pi^{*}({\bm{x}},{\bm{z}})$ as $n\rightarrow\infty$ (De Bortoli et al., 2021), and this procedure is commonly referred to as iterative proportional fitting (IPF) (Fortet, 1940; Kullback, 1968; Ruschendorf, 1995; Reich, 2019) or Sinkhorn updates (Cuturi, 2013). IPF can straightforwardly be modified to the path space setting of (23), and the resulting updates coincide with the Föllmer drift updates discussed in Section C.3, see (Vargas et al., 2021a). In applications, however, IPF is faced with the following challenges:
@@
    The reference distribution $r({\bm{x}},{\bm{z}})$ (or the reference vector field $f_{t}$) enters the iterations in (4.1) only through the initialisation. As a consequence, numerical errors accumulate, and it is often observed that the Schrödinger prior is ‘forgotten’ as IPF proceeds (Vargas et al., 2021a; Fernandes et al., 2021; Shi et al., 2023).
To address these issues, we first establish a connection to expectation-maximization (EM) (Dempster et al., 1977), originally devised for finding maximum likelihood estimates in models with latent (or hidden) variables. According to Neal and Hinton (1998), the EM-algorithm can be described in the setting from the introduction, and written in the form
$\theta_{n+1}=\operatorname*{arg\,min}_{\theta}\mathcal{L}_{D_{\mathrm{KL}}}(\phi_{n},\theta),$
$\phi_{n+1}=\operatorname*{arg\,min}_{\phi}\mathcal{L}_{D_{\mathrm{KL}}}(\phi,\theta_{n+1}),$ (25)
with $\mathcal{L}_{D_{\mathrm{KL}}}$ defined as in (5). If the initialisations are matched appropriately, the following result establishes an exact correspondence between the IPF updates in (4.1) and the EM updates in (4.1):
Proposition 4.1 (EM $\iff$ IPF).
Assume that the transition densities $p^{\theta}({\bm{x}}|{\bm{z}})$ and $q^{\phi}({\bm{z}}|{\bm{x}})$ are parameterised with perfect flexibility, [^6] In precise terms, we assume that for any transition density $p({\bm{x}}|{\bm{z}})$, there exists $\theta_{*}$ such that $p({\bm{x}}|{\bm{z}})=p^{\theta_{*}}({\bm{x}}|{\bm{z}})$. Likewise, we assume that for any transition density $q({\bm{z}}|{\bm{x}})$, there exists $\phi_{*}$ such that $q({\bm{z}}|{\bm{x}})=q^{\phi_{*}}({\bm{z}}|{\bm{x}})$. and furthermore that the EM-scheme (4.1) is initialised at $\phi_{0}$ in such a way that $q^{\phi_{0}}({\bm{z}}|{\bm{x}})=r({\bm{z}}|{\bm{x}})$. Then the IPF iterations in (4.1) agree with the EM iterations in (4.1) for all $n\geq 1$, in the sense that
[…]
@@
See Appendix E. ∎
Corollary 4.2 (Path space EM).

### v6

@@ E.1.1 Discretisation and connection to DNFs (diffusion normalising flows)
The time discretised RND in equation (E.1.1) can be expressed as the ratio of the transition densities corresponding to two discrete-time Markov chains $\mu({\bm{y}}_{0})q^{a}({\bm{y}}_{1:K}|{\bm{y}}_{0})/p^{b}({\bm{y}}_{0:K-1}|{\bm{y}}_{K})\nu({\bm{y}}_{K})$ with ${\bm{y}}_{0:K}\sim q^{a}({\bm{y}}_{1:K}|{\bm{y}}_{0})\mu({\bm{y}}_{0})$. As a result considering $\nu(x)=\hat{\nu}({\bm{x}})/Z$ and the IS estimator $\hat{Z}=p^{b}({\bm{y}}_{0:K-1}|{\bm{y}}_{K})\hat{\nu}({\bm{y}}_{K})/\mu({\bm{y}}_{0})q^{a}({\bm{y}}_{1:K}|{\bm{y}}_{0})$ it follows that $\mathbb{E}_{q^{a}({\bm{y}}_{1:K}|{\bm{y}}_{0})\mu({\bm{y}}_{0})}[\ln\hat{Z}]$ is an ELBO of $\hat{Z}$ (e.g. $\mathbb{E}_{q^{a}({\bm{y}}_{1:K}|{\bm{y}}_{0})\mu({\bm{y}}_{0})}[\ln\hat{Z}]\leq\ln{Z}$).
Whilst superficially simple, Remark 10 guarantees that normalizing constant estimators arising from our discretisation do not overestimate the true normalizing constant. This result is beneficial in practice as it allows us to compare estimators possessing this property by selecting the one with the largest value. As highlighted in Vargas et al. (2023a) many SDE discretisations can result in estimators that do not yield an ELBO: for example, the estimators used in Berner et al. (2022) can result in overestimating the normalising constant. Note similar remarks have been established in the context of free energy computation and the Jarzynski equality see Stoltz et al. (2010, Remark 4.5).
@@
E.2 Proof of Proposition 3.3 (Controlled Crooks’ fluctuation theorem and the Jarzinky equality)
Proof.
Following the computations in Appendix D.1 and using the formulae (29), we compute
@@
$\operatorname{ln}\left(\frac{\mathrm{d}\overrightarrow{\mathbb{P}}^{\mu,\sigma^{2}\nabla\ln{{\pi}}+\nabla\phi}}{\mathrm{d}\overleftarrow{\mathbb{P}}^{\nu,-\sigma^{2}\nabla\ln{{\pi}}+\nabla\phi}}\right)({\bm{Y}})=\operatorname{ln}\mu({\bm{Y}}_{0})-\operatorname{ln}\nu({\bm{Y}}_{T})$
$+\int_{0}^{T}\nabla\operatorname{ln}{\pi}_{t}({\bm{Y}}_{t})\circ\mathrm{d}{\bm{Y}}_{t}-\tfrac{1}{\sigma^{2}}\int_{0}^{T}\nabla\phi_{t}({\bm{Y}}_{t})\cdot\nabla\operatorname{ln}{\pi}_{t}({\bm{Y}}_{t})\,\mathrm{d}t-\int_{0}^{T}\Delta\phi({\bm{Y}}_{t})\,\mathrm{d}t.$
Then via Itô’s lemma applied to the unnormalised annealed log target $\ln\hat{\pi}_{t}=\ln{\pi}_{t}-\ln{\mathcal{Z}}_{t}$ we have
$\ln\hat{\pi}_{T}({\bm{Y}}_{T})$ $-\ln\hat{\pi}_{0}({\bm{Y}}_{0})-\int_{0}^{T}\partial_{t}\ln\hat{\pi}_{t}({\bm{Y}}_{t})\,\mathrm{d}t=\int_{0}^{T}\nabla\operatorname{ln}\hat{\pi}_{t}({\bm{Y}}_{t})\circ\mathrm{d}{\bm{Y}}_{t}=\int_{0}^{T}\nabla\operatorname{ln}\pi_{t}({\bm{Y}}_{t})\circ\mathrm{d}{\bm{Y}}_{t},$
thus we arrive at
@@
$\operatorname{ln}\left(\frac{\mathrm{d}\overrightarrow{\mathbb{P}}^{\mu,\sigma^{2}\nabla\ln{{\pi}}+\nabla\phi}}{\mathrm{d}\overleftarrow{\mathbb{P}}^{\nu,-\sigma^{2}\nabla\ln{{\pi}}+\nabla\phi}}\right)({\bm{Y}})=\operatorname{ln}\mu({\bm{Y}}_{0})-\operatorname{ln}\nu({\bm{Y}}_{T})+\ln\hat{\pi}_{T}({\bm{Y}}_{T})-\ln\hat{\pi}_{0}({\bm{Y}}_{0})$
$-\int_{0}^{T}\partial_{t}\ln\hat{\pi}_{t}({\bm{Y}}_{t})\,\mathrm{d}t-\tfrac{1}{\sigma^{2}}\int_{0}^{T}\nabla\phi_{t}({\bm{Y}}_{t})\cdot\nabla\operatorname{ln}\pi_{t}({\bm{Y}}_{t})\,\mathrm{d}t-\int_{0}^{T}\Delta\phi_{t}({\bm{Y}}_{t})\,\mathrm{d}t.,$
for arbitrary initial and final densities $\mu$ and $\nu$. Crooks’ generalised fluctuation theorem (Crooks, 1999) now follows from taking $\phi=0$, and the controlled version in Proposition 3.3 follows from $\mu=\pi_{0}$ and $\nu=\pi_{T}$. Finally notice that:
@@
11 11 $=\mathbb{E}_{\overrightarrow{\mathbb{P}}^{\mu,\sigma^{2}\nabla\ln{{\pi}}}}\left[\left(\frac{\mathrm{d}\overrightarrow{\mathbb{P}}^{\mu,\sigma^{2}\nabla\ln{{\pi}}}}{\mathrm{d}\overleftarrow{\mathbb{P}}^{\nu,-\sigma^{2}\nabla\ln{{\pi}}}}\right)^{-1}\right]$
@@
$=\mathbb{E}_{\overrightarrow{\mathbb{P}}^{\mu,\sigma^{2}\nabla\ln{{\pi}}}}\left[\exp\left(-\operatorname{ln}\mu({\bm{Y}}_{0})+\operatorname{ln}\nu({\bm{Y}}_{T})-\ln\hat{\pi}_{T}({\bm{Y}}_{T})+\ln\hat{\pi}_{0}({\bm{Y}}_{0})+\int_{0}^{T}\partial_{t}\ln\hat{\pi}_{t}({\bm{Y}}_{t})\,\mathrm{d}t\right)\right],$
$\mathbb{E}_{\overrightarrow{\mathbb{P}}^{{\pi}_{0},\sigma^{2}\nabla\ln{{\pi}}}}\left[\exp\left(\int_{0}^{T}\partial_{t}\ln\hat{\pi}_{t}({\bm{Y}}_{t})\,\mathrm{d}t\right)\right]=e^{-(\ln{\mathcal{Z}}_{0}-\ln{\mathcal{Z}}_{T})}.$
We want to highlight that in (Vaikuntanathan & Jarzynski, 2008, Equations 10-14) ; we can see a similar formulation to our proposed generalised Crooks’ fluctuation theorem, that said Vaikuntanathan & Jarzynski (2008) seems to pose this as a conjecture providing no rigorous proof. Furthermore, unlike our work, they do not formulate this result through SDEs, which we believe we are the first to do. In short, our work and concurrently Zhong et al. (2023) are the first to provide a rigorous treatment in establishing the escorted version of Crooks’ fluctuation theorem.
E.3 Proof of Proposition 3.1: EM $\iff$ IPF
In applications, IPF is faced with the following challenges:
@@
    The reference distribution $r({\bm{x}},{\bm{z}})$ (or the reference vector field $f_{t}$) enters the iterations in (18) only through the initialisation. As a consequence, numerical errors accumulate, and it is often observed that the Schrödinger prior is ‘forgotten’ as IPF proceeds (Vargas et al., 2021a; Fernandes et al., 2021; Shi et al., 2023).
Thus to address these challenges this section will focus on establishing the connection between EM and IPF which in turn will provide us with a family of algorithms that circumvent the sequential nature of IPF, further bridging variational inference and entropic optimal transport.
@@
The proof proceeds by induction.
To begin with, the update formula in (18a) implies that
The induction step is split (depending on whether $n$ is odd or even):
1.) First assume that the first line of (20) holds for a fixed odd $n\geq 1$. Our aim is to show that this implies that
$\pi^{n+1}({\bm{x}},{\bm{z}})=p^{\theta_{(n+1)/2}}({\bm{x}}|{\bm{z}})\nu({\bm{z}}),$ (71)
that is, the second line of (20) with $n$ replaced by $n+1$. From (18b), we see that
Again, we enforce the marginal constraint by setting $\pi({\bm{x}},{\bm{z}})=\pi({\bm{z}}|{\bm{x}})\nu({\bm{z}})$ and proceed as in (70) to obtain $\pi^{n+1}({\bm{x}},{\bm{z}})=\pi^{n}({\bm{x}}|{\bm{z}})\nu({\bm{z}})$. The statement in (71) is therefore equivalent to $\pi^{n}({\bm{x}}|{\bm{z}})=p^{\theta_{(n+1)/2}}({\bm{x}}|{\bm{z}})$. To show this, we observe from the EM-scheme in (19) that
In combination with the second line of (20) and the definition of $\mathcal{L}_{D}(\phi,\theta)$ in (5), we obtain
2.) Assume now that the second line of (20) holds for a fixed even $n\geq 2$. We need to show that the first line holds with $n$ replaced by $n+1$, that is,
$\pi^{n+1}({\bm{x}},{\bm{z}})=q^{\phi_{n/2}}({\bm{z}}|{\bm{x}})\mu({\bm{x}}).$
Remark 11 (Extension to $f$-divergences).
The proof does not make use of specific properties of $D_{\mathrm{KL}}$, other than that it satisfies the chain rule. As a consequence, the statement of Proposition 3.1 straightforwardly extends to other divergences with this property, in particular to $f$-divergences, see Proposition 6 in (Baudoin, 2002).
E.4 Drift based EM
As remarked in the previous subsection, the proof of the equivalence between IPF and EM in path space follows the exact same lines, replacing the chain rule of $D_{\mathrm{KL}}$ with the (slightly more general) disintegration theorem (Léonard, 2014b). In this section, we provide a direct extension to the control setting, yielding yet another IPF-type algorithm and motivating certain design choices for the family of methods we study.
Corollary E.2 (Path space EM).